A bullet fired at an angle θ = 60degress with a velocity of 30 m/s.



Acceleration due to gravity is 10 m/s2. What is the time interval to



reach the maximum height?

1
Expert's answer
2022-05-10T13:20:55-0400
"\Delta t=\frac{v_0\sin\theta}{g}=\frac{30\sin60^\circ}{10}=2.6\;\rm s"